OpenStudy (anonymous):

If 2y-5x=5 what is the smallest integral value of x for which y>100?

OpenStudy (anonymous):

anyone please solve this

OpenStudy (anonymous):

just try plugging in y=100 then solve.

OpenStudy (anonymous):

ok, let me try

OpenStudy (anonymous):

x=39

OpenStudy (anonymous):

but the answer is 20

OpenStudy (tkhunny):

Let's explore and see what is REALLY going on. 2y - 5x = 5 x = (2y - 5)/5 x = (2(100) - 5)/5 = (200 - 5)/5 = 195/5 = 39 y = (5 + 5x)/2 y = (5 + 5(38))/2 = (5 + 190)/2 = 195/2 = 97.5 y = (5 + 5(39))/2 = (5 + 195)/2 = 200/2 = 100 y = (5 + 5(40))/2 = (5 + 200)/2 = 205/2 = 102.5 The correct answer, the the problem posed, is x = 40 If it had said \(y \ge 100\), the answer would be x = 39. "20" is simply incorrect.

OpenStudy (anonymous):

ahh thanks tkhunny